OpenStudy (emilyjones284):

Simplify the expression. 15^-1 x^2 y^-3 z^-3/ 3^-2 x^-5 y^5

KjSaif (kjsaif):

Well this is going to hurt my head

OpenStudy (kalebj):

-35x8y13z11

OpenStudy (kalebj):

Equation at the end of step 1 : (((3x•(y2))•(z4))2)•((0-((3x2•y3)•z))3) Step 2 : Equation at the end of step 2 : ((3xy2 • (z4))2) • -33x6y9z3 Step 3 : Multiplying exponential expressions : 3.1 x2 multiplied by x6 = x(2 + 6) = x8 Multiplying exponential expressions : 3.2 y4 multiplied by y9 = y(4 + 9) = y13 Multiplying exponential expressions : 3.3 z8 multiplied by z3 = z(8 + 3) = z11 Multiplying exponents : 3.4 32 multiplied by 33 = 3(2 + 3) = 35 Final result : -35x8y13z11

KjSaif (kjsaif):

\[3y^2 / 5x^3z^3\]